# A Dual-Motor Actuator for Ceiling Robots with High Force and High Speed Capabilities

### v1

@@ Abstract
Patient transfer devices allow to move patients passively in hospitals and care centers. Instead of hoisting the patient, it would be beneficial in some cases to assist their movement, enabling them to move by themselves. However, patient assistance requires devices capable of precisely controlling output forces at significantly higher speeds than those used for patient transfers alone, and a single motor solution would be over-sized and show poor efficiency to do both functions. This paper presents a dual-motor actuator and control schemes adapted for a patient mobility equipment that can be used to transfer patients, assist patient in their movement, and help prevent falls.
The prototype is shown to be able to lift patients weighing up to 318 kg, to assist a patient with a desired force of up to 100 kg with a precision of 7.8%. Also, a smart control scheme to manage falls is shown to be able to stop a patient who is falling by applying a desired deceleration.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I INTRODUCTION
Patient transfer devices are commonly used in medical facilities to assist in moving patients at a controlled pace by supporting their full weight. Studies suggest self-movements for patients with the device assisting them are beneficial [1] [2] instead of lifting their full weight. While assistive equipments [3] [4] [5] have been successfully utilized in rehabilitation settings, their widespread implementation in hospital rooms remains cost-prohibitive. Improving existing patient transfer equipment [6] [7] [8] with rehabilitation features can enhance patient mobility. This approach ensures continued support for patients with limited mobility while facilitating rehabilitation exercises for those capable of more independent movements.
Commercial lifting devices aim for versatility for both patient transfer and patient assistance [9], yet their maximum speed of 0.1 m/s falls short for certain daily activities. For instance, during a sit-to-stand (standing up from a sitting position), the center of mass of a patient can reach a vertical velocity of 0.35 m/s [10], making this system unable to follow the patient for all their movements. Therefore, the mechanical requirements for patient assistance differ significantly from those for patient transfer. For patient transfer, the multifunctional lift winches the full weight (high force) and uses velocity control to move the patient. For patient assistance, the multifunctional lift unloads a certain percentage of the weight of the patient while following their movements. This means working at high speed and low force, and controlling the force applied on the user. The multifunctional lift also needs to be able to prevent the patient from falling to the floor (fall prevention).
@@
Addressing these divergent requirements with a single-motor actuator presents challenges due to the inherent trade-offs between force and speed capabilities. While lightly-geared electric motors excel in low-force, high-speed applications, they prove inefficient for high-force tasks typical in patient transfer scenarios. Moreover, augmenting the gear ratio like in available transfer devices limits the maximum velocity and increases the reflected inertia and friction, penalizing the force tracking accuracy.
Variable gear ratio systems like in car powertrains would address the divergent requirements for both applications without oversizing the motor. The system could downshift to a large reduction ratio to achieve high force capabilities during patient transfers, and upshift to a small reduction ratio for high-speed capabilities while assisting a patient.
@@
However, the system needs to transition from high-speed to high-force to safely stop a patient who would fall during a training session, which is not feasible with a regular variable gear system. A dual-motor actuator can offer seamless shifting between two distinct configurations since both motors can be connected to the output at the same time. This type of actuator has been explored for a wide range of robotic applications. T. Takayama and al.[11] have used a force-magnification drive for a robotic hand that can move at high speed and apply high force when gripping an object. But the high-speed mode is not backdrivable which would penalize force tracking accuracy if used with a force controller. To overcome those problems, A. Lecavalier and al. [12] present a bimodal hydrostatic actuator for robotic legs with one strong mode when the leg is supporting the weight and one fast mode for the foot placement when the leg is swinging freely. This solution leads to a lighter and more efficient actuator [13], but the range of motion is limited by the stroke of the hydraulic cylinder. A. Girard and al. [14] have developed a dual-motor actuator with a planetary differential gearbox and locking brakes to control which motor is giving energy to the output. This offers a compact solution with an unlimited range of motion at the output [15] [16]. However, the transition between both modes needs to happen at low speed due to the brake’s limitations. This is problematic for fall prevention since the system needs to take a high load at high speed. Having a brake with controllable force output, like a disk brake, would help control the deceleration of the patient and improve their comfort.
Series elastic actuators (SEA) could also improve force control by negating motor friction and inertia [17] [18] [19], but the spring limits the displacement for the assistance mode. To control the output, a force sensor can be mounted in series with the load and use its signal for a closed-loop force control algorithm [17] [20]. A friction compensation algorithm can also be used instead to control the force without the need for an extra component [21]. Both these strategies were implemented and compared for this particular application.
In summary, although dual-motor actuators show promise for applications with discrete points of operation, their integration into healthcare lifting and assistive devices remains incomplete. Furthermore, using a disk brake to control the transition between modes has not been widely explored in this context. These areas represent opportunities for future innovation in the field of patient transfer and assistance technology.
@@
This paper presents the design and control of a novel dual-motor multifunctional ceiling-based device that extends the capabilities of current healthcare lifting devices up to the capabilities of current assistive devices, with little compromise on system mass and efficiency. Figure 1 gives an overview of the system and its operating modes. The contributions are: A) a novel control scheme for a dual-motor actuator with a disk brake for a seamless transition at high speed and high force. B) The analysis of different actuator architectures with a comparison of their performances. C) The implementation of a dual-motor actuator for medical equipment used for patient transfer and patient assistance. D) A force control friction compensation algorithm that utilizes both motors to improve the performances over a single motor solution. Section II presents the system requirements. In section III, the selected motor architecture is introduced and compared to other relevant designs. This section also describes the dynamical model and the research prototype designed. Section IV presents control algorithms for three operating modes: patient transfer, patient assistance, and fall prevention with experimental results shown in section V.
@@ II-A Patient Transfer
Patient transfer is for hoisting patients up to 272 kg (600 lb) at 0.04 m/s (107 W). This fits within the range of typical medical lifting devices, which can lift patients weighing between 200 and 450 kg at a speed between 0.03 and 0.05 m/s [6] [7] [8]. Safe and precise manipulation of the patient limits the desired speed in this mode.
@@ III Actuator Design and Modeling
This section presents in detail the proposed multifunctional lift. First, the mechanical principle is described with a simplified comparison showing how this design is superior to more typical design strategies for reaching the same functions. Then, the equations of motion are detailed. Finally, the experimental prototype is detailed with the actual specifications reached.
III-A Mechanical Principle
The two-speed architecture consists of two motors and a planetary gear train. One motor is a highly geared electric motor (EM1), while the other is a lightly geared electric motor (EM2). As shown in Figure 4, the drum output that is attached to the patient connects to the carrier of the planetary gearbox. The planetary is thus used as a differential. This results in a redundant system, the displacement of each motor adds up to drive the output.
III-B Design Comparison
@@
Figure 3 shows alternative concepts possible for a ceiling device: 1) one small highly-geared motor; 2) one powerful lightly-geared motor; 3) one motor, two gear ratios, and two clutches; and 4) two motors, two gear ratios, one brake (as proposed). For each concept, Table I gives a design proposition based on a list of commercial components and the identification of proper reduction ratios that allows reaching the force requirements for HF and HS modes. Motors are frameless models, so a factor of 1.5 is used for mass, for considering housing mass. It is assumed that by using electromagnetic disc clutches and brakes (which can slip), seamless transitions between HF and HS modes would be possible by controlling the dissipation forces. The mass is for comparison only since it excludes gearboxes, main frame, batteries, etc. Transmission efficiency is assumed 100% for simplification.
@@
The single small motor design is the lightest (and probably cheapest) design. It has limited dynamics capabilities in HS mode since the inertia and speed would hinder patient accelerations and natural movements. Still, this simple solution may be sufficient for early rehabilitation and aged patients. The single big motor design can reach all force and speed requirements by the use of a powerful motor. The reflected inertia is by comparison low which means better force control for dynamic motions. Mass (and indirectly cost) is however much higher. The selection of the motor size and ratio is a compromise between reflected inertia and motor mass. This system would potentially cause safety hazards due to its high power capabilities. Finally, the last two dual-speed designs can reach all requirements without the need for powerful motors. The fourth design has a better mass/inertia advantage because: 1) disc brakes are lighter than disc clutches (36% lighter for the MikiPulley models) and 2) for the third design, the reflected inertia of the HF clutch strongly affects the total inertia at the output for the HS mode ($R_{1}$ being large), which means there is a mass vs. inertia compromise when choosing the reduction ratio between the HF clutch and the motor. One drawback of the two-speed designs is that they require more components, especially more gear stages. The fourth design offered the best tradeoff between performance and size, and was thus chosen for the prototype.
@@ III-C Equations of Motion
III-D Prototype
A prototype of a multifunctional lift was assembled, including the dual-motor actuator and a drum that winds a strap as pictured in Figure 5. The prototype includes a support for a harness that will be installed on a patient. For EM1, a Maxon RE40 150 W brushed motor with a 66:1 integrated gearbox is coupled to the ring gear of the planetary gearbox. For EM2, a Tecnotion QTR-A-78-25 is coupled directly to the sun gear. The custom gearbox includes a planetary gear stage and other spur gear stages. The total reduction ratios are respectively $R_{1}=600$ and $R_{2}=18$ for EM1 and EM2. The output of the custom gearbox is connected to a drum (average 0.04 m radius $r$). Encoders are implemented on both motors, and a load cell measures the load applied on the strap. EM1 is powered by an Escon50/5 driver and EM2 with an ODriveV3.6 driver. Both motor drives allow for current and velocity control at low level. A disc brake is connected to EM2 to stop it from being backdriven by EM1 during HF mode and for fall prevention. The disc brake is made of carbon fiber to minimize the reflected inertia in HS mode. The spring is in series between the servomotor (DS3235SG, 35 kg) and the brake caliper (series-elastic actuation) so that the position of the servomotor roughly controls the braking torque. An Arduino Mega is used for the software control with a 250 Hz control loop.
The prototype’s theoretical capabilities are outlined in Table II. In HF mode, the load is limited by the 66:1 gearbox rated torque. In HS mode, the peak force is reached when sending twice the nominal current of the motor. All specifications for speed and force in HF and HS modes were reached. The calculated reflected inertia at output includes the motor frameless rotor (59%), the custom shaft (22%), and the disc brake (19%). The lightly-geared design is thus expected to properly track sit-to-stands and normal walking. It also leads to a reasonable reflected inertia (as opposed to the highly-geared line) relative to the nominal assistive forces and patient’s weight.
Prototype Guldmann Aretech
Mode HF HS HF HS HF HS
Ratio 600:1 18:1 N/A N/A N/A N/A
Max force 318 59 400 100 204 91
(kgf) (100 peak)
Loaded 0.05 0.55 0.06 0.1 N/A N/A
velocity (m/s) (34 @ peak)
Reflected 3427 8.5 N/A N/A N/A N/A
inertia (kg)
@@ IV CONTROL ALGORITHMS
HF and HS mode transitions occur by opening or closing the brake. If a fall is detected in HS mode, the brake and EM2 work simultaneously to decelerate the patient before downshifting. To help downshift faster, EM1 can accelerate downward to slow down EM2 (see equation 8). Once the velocity of EM2 is zero, the brake closes to its maximal braking force and the system becomes in HF mode. To help the patient recover and continue his activity, the system then lifts him to the position where the fall was detected. In the following section, controller details are presented for each operating mode.
@@ IV-B Patient Assistance
For daily activities assistance, the lift unloads weight to the patient while following their movements. This requires the multifunctional lift to be in HS mode and to control the force at the output rather than the position or velocity as in transfer mode. The force control algorithms that were tested are presented in Figure 7: a) an open-loop current control, b) a friction compensation algorithm, c) a friction compensation algorithm with an offset on the speed by leveraging the nullspace, d) a PID current controller and e) a disturbance observer (DOB).
a) The motor torque is obtained by knowing the motor constant ($k_{t}$) and adjusting the current of EM2. This works well for lightly geared electric motors, but the friction can induce an undesired error on the force output. This error can be measured in terms of speed (see Figure 8) and a dry and viscous friction model can be estimated using a linear equation (using the absolute value for symmetry) multiplied by a hyperbolic tangent function:
$\tau_{f}=f(\omega_{2})=(b|\omega_{2}|+c)tanh(\omega_{2})$ (9)
b) The friction estimation can then be subtracted to the desired force to compensate for the friction error and help reduce the force tracking error. One issue with friction compensation is when trying to compensate near zero speed since a more complex friction model would be necessary for considering hysteresis. c) With a dual-motor actuator, EM1 can run at a constant speed while EM2 is applying a constant force. This offsets the speed of EM2 from the output speed, which reduces the time EM2 is at zero speed (see equation 1, since the patient is often near zero speed (when they’re standing still or sitting down). d) A load cell is typically required on ceiling lifts for measuring patient’s weight, the measured force can be sent to a simple closed-loop PID controller. e) A disturbance observer was also implemented to compare a different and more elaborate control law.
@@
Figure 8: Measure of the friction in EM2 without and with speed offset from EM1.

### v2

@@ Abstract
Patient transfer devices allow to move patients passively in hospitals and care centers. Instead of hoisting the patient, it would be beneficial in some cases to assist their movement, enabling them to move by themselves and reducing hospitalization time. However, patient assistance requires devices capable of precisely controlling output forces at significantly higher speeds than those used for patient transfers alone, and a single motor solution would be over-sized and show poor efficiency to do both functions. This paper presents a ceiling robot, using a dual-motor actuator and adapted control schemes, that can be used to transfer patients, assist patients in their movement, and help prevent falls. The prototype is shown to be able to lift patients weighing up to 318 kg, and to assist a patient with a desired force of up to 100 kg with a precision of 7.8%. Also, a smart control scheme to manage falls is shown to be able to stop a patient who is falling by applying a desired deceleration.
Index Terms:
Force Control; Physically Assistive Devices; Rehabilitation Robotics; multifunctionnal actuator;
@@ I INTRODUCTION
Patient transfer devices are commonly used in medical facilities to help caregivers move patients at a controlled pace by supporting their full weight. Studies suggest self-movements for patients with the device assisting them are beneficial [1] [2] instead of lifting their full weight. This can help improve healing time and reduce rehospitalization risks [3][4]. While assistive equipment [5] [6] [7] have been successfully utilized in rehabilitation settings, their widespread implementation in hospital rooms remains cost-prohibitive. Improving existing patient transfer equipment [8] [9] [10] with rehabilitation features can enhance patient mobility. This approach ensures continued support for patients with limited mobility while facilitating rehabilitation exercises for those capable of more independent movements.
Commercial lifting devices aim for versatility for both patient transfer and patient assistance [11], yet their maximum speed of 0.1 m/s falls short for certain daily activities. For instance, during a sit-to-stand (standing up from a sitting position), the center of mass of a patient can reach a vertical velocity of 0.35 m/s [12], making this system unable to follow the patient for all their movements. Using a smaller gear ratio for the actuator can increase it’s maximum speed and enable following the patient during rehabilitation. Unfortunately, this decreases the maximum force at the output making this solution unsuited for both applications.
Therefore, the mechanical requirements for patient assistance differ significantly from those for patient transfer. For patient transfer, the lift winches the full patient’s weight (high force) and uses velocity control to move the patient. For patient assistance, the lift unloads a certain percentage of the patient’s weight while following their movements. This means working at high speed and low force, and controlling the force applied on the user. The lift also needs to be able to prevent the patient from falling to the floor (fall prevention).
@@
Addressing the divergent requirements of patient transfer and assistance with a single-motor actuator presents challenges due to the trade-offs between force and speed capabilities. While lightly-geared electric motors excel in low-force, high-speed applications, they prove inefficient for high-force tasks typical in patient transfer scenarios. Increasing the gear ratio like in available transfer devices limits the maximum velocity and increases the reflected inertia and friction, penalizing the force tracking accuracy.
Alternative approaches have been explored to tackle similar problems: 1) Oversized electric motor: One simple solution to reach both modes of operation is to use a bigger motor with enough force and speed capabilities for patient transfer and rehabilitation [7] [13]. This solution increases the actuator’s weight (and therefore costs). For a ceiling base actuator, this also means the patient needs to carry a heavier device while walking during rehabilitation. This is undesirable for patients who already need assistance while walking. 2) Force feedback: To control the output, a force sensor can be mounted in series with the load and use its signal for a closed-loop force control algorithm [7] [14] [15]. A friction compensation algorithm can also be used instead to control the force without the need for a load cell [16]. Despite using a force control algorithm, the performances are limited by the actuator used. 3) Series elastic actuator (SEA): SEA have been explored in research for patient assistance devices to improve force control by negating motor friction and inertia [7] [17] [18], but the spring limits the displacement for the assistance mode and needs to be locked before transfer mode to support the full weight of the patient. This is problematic for care centers since it adds time for patient transfer tasks. Adding a spring also doesn’t improve the maximum speed during rehabilitation which is again limited by the actuator. 4) Variable gear-ratio systems: Similarly to car powertrains, variable gear-ratios could address the divergent requirements for both applications without oversizing the motor. The system could downshift to a large reduction ratio to achieve high force capabilities during patient transfer and upshift to a small reduction ratio for high-speed capabilities while assisting a patient. However, transitioning from high-speed to high-force is crucial to safely stop a patient from falling during a training session, which is not feasible with a regular variable gear-ratio system [19].
@@
A dual-motor actuator can offer seamless shifting between two distinct configurations since both motors can be connected to the output at the same time. This type of actuator has been explored for a wide range of robotic applications. T. Takayama and al. [20] have used a force-magnification drive for a robotic hand that can move at high speed and apply high force when gripping an object. But the high-speed mode is not backdrivable which would penalize force tracking accuracy if used with a force controller. To overcome those problems, A. Lecavalier and al. [21] present a bimodal hydrostatic actuator for robotic legs with one strong mode when the leg is supporting the weight and one fast mode for the foot placement when the leg is swinging freely. This solution leads to a lighter and more efficient actuator [22], but the range of motion is limited by the stroke of the hydraulic cylinder. A. Girard and al. [19] have developed a dual-motor actuator with a planetary differential gearbox and locking brakes to control which motor is giving energy to the output. This offers a compact solution with an unlimited range of motion at the output [23]. However, the transition between both modes needs to happen at low speed due to the brake’s limitations. This is problematic for fall prevention since the system needs to take a high load at high speed. Having a brake with controllable force output, like a disk brake, would help control the deceleration of the patient and improve their comfort. In summary, although dual-motor actuators show promise for applications with discrete points of operation, applying this technology to patient handling equipment is not explored yet and includes specific technical challenges such as managing smooth transitions while fighting a large backdriving force for fall prevention. Addressing those technical challenges with a modified mechanical design and control scheme is the main contribution of this paper.
@@
This paper presents the design and control of a novel dual-motor multifunctional ceiling-based device that extends the capabilities of current healthcare lifting devices up to the capabilities of current assistive devices, with little compromise on system mass and efficiency. Figure 1 gives an overview of the system and its operating modes. The contributions are: a) a novel control scheme for a dual-motor actuator with a disk brake for a seamless transition at high speed and high force (a specific requirement for fall prevention), b) a comparative analysis of four actuator design architectures, c) a novel friction compensation algorithm that leverage the nullspace of the dual-motor actuator, and d) a full-sized, fully functional prototype of a multi-function ceiling lift using the proposed dual-motor actuator technology. Section II presents the system requirements. In section III, the selected motor architecture is introduced and compared to other relevant designs. This section also describes the dynamical model and the research prototype designed. Section IV presents control algorithms for three operating modes: patient transfer, patient assistance, and fall prevention with experimental results shown in section V.
@@ II-A Patient Transfer
Patient transfer is for hoisting patients up to 272 kg (600 lb) at 0.04 m/s (107 W). This fits within the range of typical medical lifting devices, which can lift patients weighing between 250 and 450 kg at a speed between 0.03 and 0.05 m/s [8] [9] [10]. Safe and precise manipulation of the patient limits the desired speed in this mode.
@@ III Actuator Design and Modeling
This section presents in detail the proposed multifunctional lift. First, different generic design choices are compared to show how the chosen design is superior to more typical design strategies for reaching the same functions. Then, the final actuator design is described, including the experimental prototype and the actual specifications reached. Finally, the equations of motion are detailed.
III-A Preliminary Design Exploration
@@
Figure 3 shows alternative concepts possible for a ceiling device based on already explored actuator solutions: 1) one small highly-geared motor [11]; 2) one powerful lightly-geared motor [5]; 3) one motor, two gear ratios, and two clutches [27]; and 4) two motors, two gear ratios, one brake (as proposed). For each concept, Table I gives a design proposition based on a list of commercial components and the identification of proper reduction ratios that allows reaching the force requirements for HF and HS modes. Motors are frameless models, so a factor of 1.5 is used for mass, for considering housing mass. It is assumed that by using electromagnetic disk clutches and brakes (which can slip), seamless transitions between HF and HS modes would be possible by controlling the dissipation forces. The mass is for comparison only since it excludes gearboxes, main frame, batteries, etc. Transmission efficiency is assumed 100% for simplification.
@@
The single small motor design is the lightest (and probably cheapest) design. It has limited dynamics capabilities in HS mode since the inertia and speed would hinder patient accelerations and natural movements. Still, this simple solution may be sufficient for early rehabilitation and aged patients. The single big motor design can reach all force and speed requirements by the use of a powerful motor. The reflected inertia is by comparison low which means better force control for dynamic motions. Mass (and indirectly cost) is however much higher. The selection of the motor size and ratio is a compromise between reflected inertia and motor mass. This system would potentially cause safety hazards due to its high power capabilities. Finally, the last two dual-speed designs can reach all requirements without the need for powerful motors. The fourth design has a better mass/inertia advantage because: 1) disk brakes are lighter than disk clutches (36% lighter for the MikiPulley models) and 2) for the third design, the reflected inertia of the HF clutch strongly affects the total inertia at the output for the HS mode ($R_{1}$ being large), which means there is a mass vs. inertia compromise when choosing the reduction ratio between the HF clutch and the motor. One drawback of the two-speed designs is that they require more components, especially more gear stages. The fourth design offered the best tradeoff between performance and size, and was thus chosen for the prototype.
@@
III-B Final Design Choice and Prototype
The two-speed architecture chosen consists of two motors and a planetary gear train. One motor is a highly geared electric motor (EM1), while the other is a lightly geared electric motor (EM2). As shown in Figure 5, the drum output that is attached to the patient connects to the carrier of the planetary gearbox. The planetary is thus used as a differential. This results in a redundant system, the displacement of each motor adds up to drive the output.
A prototype of a multifunctional lift was assembled, including the dual-motor actuator and a drum that winds a strap as pictured in Figure 4. The prototype includes a support for a harness that will be installed on a patient. For EM1, a Maxon RE40 150 W brushed motor with a 66:1 integrated gearbox is coupled to the ring gear of the planetary gearbox. For EM2, a Tecnotion QTR-A-78-25 is coupled directly to the sun gear. The custom gearbox includes a planetary gear stage and other spur gear stages. The total reduction ratios are respectively $R_{1}=600$ and $R_{2}=18$ for EM1 and EM2. The output of the custom gearbox is connected to a drum (average 0.04 m radius $r$). Encoders are implemented on both motors, and a load cell measures the load applied on the strap. EM1 is powered by an Escon50/5 driver and EM2 with an ODriveV3.6 driver. Both motor drives allow for current and velocity control at low level. Since EM1 is much stronger than EM2, a disk brake is connected to EM2 to stop it from being backdriven by EM1 during HF mode and for fall prevention. When the disk is in braking position, it locks the sun gear of the gearbox. The disk brake is made of carbon fiber to minimize the reflected inertia in HS mode. The spring is in series between the servomotor (DS3235SG, 35 kg) and the brake caliper (series-elastic actuation) so that the position of the servomotor roughly controls the braking torque. An Arduino Mega is used for the software control with a 250 Hz control loop.
The prototype’s theoretical capabilities are outlined in Table II. In HF mode, the load is limited by the 66:1 gearbox rated torque. In HS mode, the peak force is reached when sending twice the nominal current of the motor. All specifications for speed and force in HF and HS modes were reached. The calculated reflected inertia at output includes the motor frameless rotor (59%), the custom shaft (22%), and the disk brake (19%). The lightly-geared design is thus expected to track sit-to-stands and normal walking properly. It also leads to a reasonable reflected inertia (as opposed to the highly-geared line) relative to the nominal assistive forces and patient’s weight, and the measured backdriving force is 3.2 kg.
Prototype Guldmann Aretech
Mode HF HS HF HS HF HS
Ratio 600:1 18:1 N/A N/A N/A N/A
Max force 318 59 400 100 136 68
(kgf) (100 peak)
Loaded 0.05 0.55 0.06 0.1 N/A N/A
velocity (m/s) (0.34@peak)
Reflected 3427 5.1 N/A N/A N/A N/A
inertia (kg)
@@ III-C Equations of Motion
TABLE III: Design parameters summary for both motors and the output
Effort Velocity Damping Inertia Ratio
EM1 $\tau_{1}$ $\omega_{1}$ $b_{1}$ $I_{1}$ $R_{1}$
EM2 $\tau_{2}$ $\omega_{2}$ $b_{2}$ $I_{2}$ $R_{2}$
Output $F_{m}$ $v_{0}$ $b_{0}$ $m$ $r$
@@ IV CONTROL ALGORITHMS
HF and HS mode transitions occur by opening or closing the brake. If a fall is detected in HS mode, the brake and EM2 work simultaneously to decelerate the patient before downshifting. To help downshift faster, EM1 can accelerate downward to slow down EM2 (see equation 8). Once the velocity of EM2 is zero, the brake closes to its maximal braking force and the system becomes in HF mode. To help the patient recover and continue his activity, the system then lifts them to the position where the fall was detected. In the following section, controller details are presented for each operating mode.
@@ IV-B Patient Assistance
For daily activities assistance, the lift unloads weight to the patient while following their movements. This requires the multifunctional lift to be in HS mode and to control the force at the output rather than the position or velocity as in transfer mode. There are different ways to control the output force of an actuator. a) The output torque can be obtained by knowing the motor constant ($k_{t}$) and adjusting the current of EM2. This works well for quasi direct drive motors [28], it results in a simpler controller and removes the need for a load cell. However, since EM2 is a lightly geared motor, the friction is augmented by the gearbox and can induce an undesired error on the force output. This error can be measured in terms of speed (see Figure 8) and a dry and viscous friction model [29] can be estimated using a linear equation (using the absolute value for symmetry) multiplied by a hyperbolic tangent function:
$\tau_{f}=f(\omega_{2})=(b|\omega_{2}|+c+dF_{d})tanh(\omega_{2})$ (9)
Where b is the viscous friction coefficient, c is the dry friction offset and d a scaling parameter to offset the dry friction with the desired force.
b) The friction estimation can then be subtracted to the desired force to compensate for the friction error and help reduce the force tracking error. One issue with friction compensation is when trying to compensate near zero speed since a more complex friction model would be necessary for considering hysteresis. c) With a dual-motor actuator, EM1 can run at a constant speed while EM2 is applying a constant force. This offsets the speed of EM2 from the output speed, which reduces the time EM2 is at zero speed (see equation 1), since the patient is often near zero speed (when they’re standing still or sitting). Research prototypes for assistance devices often use the feedback from a load cell to measure the force applied on the user and send it to a closed loop controller, for instance d) a PID [7] or e) a disturbance observer [15].
The force control algorithms that were tested are presented in Figure 7: a) an open-loop current control, b) a friction compensation algorithm, c) a friction compensation algorithm with an offset on the speed by leveraging the nullspace, d) a PID current controller and e) a disturbance observer (DOB).
@@
Figure 8: Measure of the friction in EM2 without and with speed offset from EM1 while a 150 N load is applied.